OpenStudy (anonymous):

Please help me with this problem: What rate of interest compounded continuously is required to double an investment in 5 years? (Round your answer to one decimal place.)

OpenStudy (anonymous):

We have the sequence x1, x1*q, x1*q^2, x1^q3, x1^q4,x1*q^5 and q^5=2. q = \[\sqrt[5]{2}\] = 1.148. q represents the (principal + interest) / principal and so to find the interest/principal (interest rate) we take 1.148 - 1 = 0.148 And so the interest rate should be 14.8%